# Improved Regret Bounds for Gaussian Process Upper Confidence Bound in Bayesian Optimization

### v2

@@ 3 Improved Regret Bound for GP-UCB
Suppose Assumptions 1 and 2. Set $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Furthermore, assume that $d,\nu,\ell$, $r$ , and $\sigma^{2}$ are fixed constants. Fix any $\delta_{\mathrm{GP}}\in(0,1)$, and set the confidence width parameter $\beta_{t}$ of GP-UCB as defined in Eq. (6) with any fixed $\delta\in(0,1-\delta_{\mathrm{GP}})$. Then, with probability at least $1-\delta_{\mathrm{GP}}-\delta$, the cumulative regret of GP-UCB (Algorithm 1) satisfies [^2] Correction for $k=k_{\mathrm{Mat\acute{e}rn}}$: In the earlier version of the paper, we missed and omitted some steps in the original proof of the lenient regret by Cai et al. [8]. Since this omission necessitates an additional assumption: $4\nu^{2}+2\nu d\geq d^{2}$ in the earlier version of the paper, it has been corrected in the present version. We are grateful to Jonathan Scarlett for pointing this out.  [^3] Correction for $k=k_{\mathrm{SE}}$: In the earlier version of the paper, we obtained $O(\sqrt{T\ln^{4}T})$ upper bound for $k=k_{\mathrm{SE}}$; however, we can reduce additional multiplicative $O(\ln T)$ term by trivial modification of the proof in Lemma 4, and we correct it in this version. We are grateful to Shion Takeno for pointing this out.
@@
The hidden constants in the above expressions may depend on $\ln(1/\delta),d,\nu,\ell,r,\sigma^{2}$, and the constants $c_{\mathrm{sup}},c_{\mathrm{gap}}$, $\rho_{\mathrm{lin}},\rho_{\mathrm{quad}},c_{\mathrm{lin}},c_{\mathrm{quad}}$ corresponding with $\delta_{\mathrm{GP}}$, which are guaranteed to exist by Lemma 2.
We would like to note the following two aspects of our results. First, the constants associated with the sample path properties defined in Lemma 2 are used solely for analyzing the regret. On the other hand, the existing algorithm provided by Scarlett [43], which shows the same $\widetilde{O}(\sqrt{T})$ regret as ours, requires prior information about these constants for the algorithm run. This is often unrealistic in practice. Secondly, our result does not imply the upper bound of Bayesian expected regret $\mathbb{E}[R_{T}]$. The main issue is that the dependence of the constants in Lemma 2 on $\delta_{\mathrm{GP}}$ is not explicitly known. We leave future work to break this limitation; however, note that the same limitation exists in the algorithm provided by Scarlett [43].
@@ 3.1 Intuitive Explanation of our Analysis
  • Case II: $I(\mathbf{X}_{T})\leq\gamma_{T}(\mathcal{X})$ is tight: Let us assume that $(\bm{x}_{t})$ is the same as the input sequence generated by the maximum variance reduction (MVR) algorithm (namely, $\forall t\in[T],\bm{x}_{t}\in\operatornamewithlimits{argmax}_{\bm{x}\in\mathcal{X}}\sigma(\bm{x};\mathbf{X}_{t-1})$[48, 52]. Then, from the discussion in Sections 2 and 5 in [48], we already know that $\gamma_{T}(\mathcal{X})\leq(1-1/e)^{-1}I(\mathbf{X}_{T})$. This suggests that $I(\mathbf{X}_{T})\leq\gamma_{T}(\mathcal{X})$ is tight up to a constant factor when $\mathbf{X}_{T}$ is realized by MVR.
@@
From Case I, we observe that $I(\mathbf{X}_{T})$ satisfies $\Theta(\ln T)\leq I(\mathbf{X}_{T})\leq\gamma_{T}(\mathcal{X})$ depending on $\mathbf{X}_{T}$. Furthermore, by comparing the input sequences in cases I and II, we expect that $I(\mathbf{X}_{T})$ becomes small if $\mathbf{X}_{T}$ concentrates around the neighborhood of $\bm{x}^{\ast}$, while $I(\mathbf{X}_{T})$ becomes large if $\mathbf{X}_{T}$ spread over entire input domain $\mathcal{X}$. Then, from the fact that the worst-case regret of GP-UCB increases sub-linearly with the speed of $O(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})})$, we can deduce that the input sequence $\mathbf{X}_{T}$ of GP-UCB will eventually concentrate around the maximizer $\bm{x}^{\ast}$ if $\bm{x}^{\ast}$ is unique and $\|f\|_{\infty}$ is not extremely small [^4] Specifically, if $T\|f\|_{\infty}\leq O(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})})$, we cannot make any claims about $\mathbf{X}_{T}$ based on the worst-case bound since any sequence $\mathbf{X}_{T}$ satisfies the worst-case bound without concentrating around maximizer. This is why our analysis technique does not improve the worst-case regret in the frequentist setting. Indeed, in the proof of the worst-case lower bound for the frequentist setting [44], the existence of the function $f$ with $T\|f\|_{\infty}=O(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})})$ is guaranteed. . We provide an illustrative image in Figure 1. Our proof is designed so as to capture the above intuition that $I(\mathbf{X}_{T})$ could be improved from $\gamma_{T}(\mathcal{X})$ to $\Theta(\ln T)$ under ”favorable” sample path $f$.
@@
Figure 1: The behavior of the realized input sequence $\mathbf{X}_{T}$ (left) and the corresponding information gain $I(\mathbf{X}_{T})$ (right) in the 1-dimensional BO problem with $\sigma^{2}=3$. The top left figure shows the objective function $f$ realized by GP under $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu=5/2$. The bottom left figure shows the histogram of the realized inputs: $(\bm{x}_{t})_{t\in[200]}$ with GP-UCB (blue) and MVR (orange) under $f$ in the top left figure. Furthermore, the right plot shows the corresponding information gain $I(\mathbf{X}_{t})$ under GP-UCB or MVR. We also plot $I(\mathbf{X}_{t})\coloneqq 0.5\ln(1+\sigma^{-2}t)$, corresponding to Case I described in Section 3.1. We can observe that the inputs selected by GP-UCB are concentrated around the maximizer from the left figure. Then, from the right figure, we also observe that the corresponding information gain increases more slowly than that of MVR, and behaves similarly to Case I on $t\geq 30$.
@@ 3.2 Proof of Theorem 3
Let $\mathcal{A}$ be an event such that the two high-probability events of the original GP-UCB proof (described in the last paragraph in Section 2) and Lemma 2 with the confidence level $\delta_{\mathrm{GP}}$ simultaneously hold. Note that event $\mathcal{A}$ occurs with probability at least $1-\delta_{\mathrm{GP}}-\delta$ from the union bound. Therefore, it is enough to prove our upper bound under $\mathcal{A}$. To encode the high-level idea in the previous section, we need to capture the concentration behavior of input sequence $\mathbf{X}_{T}$ around maximizer $\bm{x}^{\ast}$. From this motivation, given some constant $\varepsilon>0$, we decompose the regret as $R_{T}=R_{T}^{(1)}(\varepsilon)+R_{T}^{(2)}(\varepsilon)$, where:
@@
We set $\mathcal{T}(\varepsilon)=\{t\in[T]\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})>\varepsilon\}$ and $\mathcal{T}^{c}(\varepsilon)=[T]\setminus\mathcal{T}(\varepsilon)$ in the above definition. A key observation is that, if we set sufficiently small $\varepsilon$ depending on the constants in Lemma 2, the inputs $(\bm{x}_{t})$ in $R_{T}^{(2)}(\varepsilon)$ (namely, inputs $(\bm{x}_{t})$ such that $f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon$ holds) are on the locally linear or quadratic region around the maximizer $\bm{x}^{\ast}$ due to conditions 1, 3, and 4 in Lemma 2. The formal descriptions are provided in Lemma 20 in Appendix C. This fact is originally leveraged in [43] to analyze the successive elimination-based algorithm. In the analysis of GP-UCB, it enables us to analyze the behavior of the sub-input sequence $\{\bm{x}_{t}\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon\}$ through the regularity constants $c_{\mathrm{lin}}$ and $c_{\mathrm{quad}}$. Below, we formally give the upper bound for $R_{T}^{(2)}(\varepsilon)$.
@@
Suppose $(\bm{x}_{t})_{t\in[T]}$ is the input query sequence realized by the GP-UCB algorithm. Furthermore, let $\overline{\gamma}_{t}$ is the upper bound of MIG $\gamma_{t}(\mathcal{X})$ such that $\overline{\gamma}_{t}/t$ is non-increasing on $[\overline{T},\infty)$ with some $\overline{T}\in\mathbb{N}_{+}$ [^5] Namely, $\forall t\geq\overline{T},\forall\epsilon\geq 0,\overline{\gamma}_{t}/t\geq\overline{\gamma}_{t+\epsilon}/(t+\epsilon)$ and $\forall t\geq\overline{T},\gamma_{t}(\mathcal{X})\leq\overline{\gamma}_{t}$ hold for some $\overline{T}\in\mathbb{N}_{+}$.. Then, under event $\mathcal{A}$, we have
$R_{T}^{(2)}(\varepsilon)$ $\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)$
$~{}~{}~{}+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\max_{i\in[\overline{i}]}\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}},$
where $C=2/\ln(1+\sigma^{-2})$, $\overline{i}=\lfloor\log_{2}\frac{T}{\overline{T}}\rfloor+1$, $\eta_{i}=\frac{2\left(2\sqrt{C\beta_{T}(T/2^{i-1})\overline{\gamma}_{T/2^{i-1}}}+\frac{\pi^{2}}{6}\right)}{(T/2^{i-1})}$, and $\varepsilon=\min\{\varepsilon_{1},\varepsilon_{2}\}>0$ with $\varepsilon_{1}=\min\{c_{\mathrm{gap}},c_{\mathrm{lin}}\rho_{\mathrm{lin}}\}$ and $\varepsilon_{2}=\min\{c_{\mathrm{gap}},c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}\}$.
@@
$R_{T}^{(2)}(\varepsilon)=\widetilde{O}\left(\max_{i}\sqrt{T\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}}\right).$ (12)
Note that $\eta_{i}$ is decreasing as the time index $T/2^{i-1}$ of MIG increases. In other words, the input domains $\mathcal{B}_{2}(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast})$ and $\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)$ of MIG shrinks as the time index $T/2^{i-1}$ increases. This property is beneficial for obtaining a tighter upper bound than that from the existing technique. For example, under $k=k_{\mathrm{Mat\acute{e}rn}}$ with $2\nu+d\leq\nu^{2}$, we can confirm that the dominant polynomial term in MIG is canceled out by the shrinking of the input domain in MIG. Namely, we can obtain the following result under $k=k_{\mathrm{Mat\acute{e}rn}}$:
$\max_{i}\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}=\widetilde{O}(1)~{}~{}(\mathrm{as}~{}~{}T\rightarrow\infty),$ (13)
@@
The full proof is given in Appendix A.2. The remaining interest is the upper bound of $R_{T}^{(1)}(\varepsilon)$. The definition of $R_{T}^{(1)}(\varepsilon)$ is the same as the lenient regret [8], which is known to be smaller than the original regret $R_{T}$ in GP-UCB. Although Cai et al. [8] studies the frequentist setting, their proof strategy is also applicable to the Bayesian setting as described in Section 3.4 in [8]. The following lemma provides the formal statement about the upper bound of $R_{T}^{(1)}(\varepsilon)$.
Lemma 6 (Upper bound of $R_{T}^{(1)}$, adaptation of the proof of Theorem 1 in [8]).

### v3

@@ 3 Improved Regret Bound for GP-UCB
Suppose Assumptions 1 and 2 hold. Set $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Furthermore, assume that $d,\nu,\ell$, $r$ , and $\sigma^{2}$ are fixed constants. Fix any $\delta_{\mathrm{GP}}\in(0,1)$, and set the confidence width parameter $\beta_{t}$ of GP-UCB as defined in Eq. (6) with any fixed $\delta\in(0,1-\delta_{\mathrm{GP}})$. Then, with probability at least $1-\delta_{\mathrm{GP}}-\delta$, the cumulative regret of GP-UCB (Algorithm 1) satisfies
@@
The hidden constants in the above expressions may depend on $\ln(1/\delta),d,\nu,\ell,r,\sigma^{2}$, and the constants $c_{\mathrm{sup}},c_{\mathrm{gap}}$, $\rho_{\mathrm{quad}},c_{\mathrm{quad}}$ corresponding with $\delta_{\mathrm{GP}}$, which are guaranteed to exist by Lemma 2.
We would like to note the following three aspects of our results. First, the constants associated with the sample path properties defined in Lemma 2 are used solely for analyzing the regret. On the other hand, the existing algorithm provided by scarlett2018tight, which shows the same $\widetilde{O}(\sqrt{T})$ regret as ours, requires prior information about these constants for the algorithm run. This is often unrealistic in practice. Secondly, our result does not imply the upper bound of Bayesian expected regret $\mathbb{E}[R_{T}]$. The main issue is that the dependence of the constants in Lemma 2 on $\delta_{\mathrm{GP}}$ is not explicitly known. We leave future work to break this limitation; however, note that the same limitation exists in the algorithm provided by scarlett2018tight. Thirdly, our results in Theorem 3 only focus on the dependence of the total step size $T$ in the regret. Therefore, we cannot claim any improvements of the regret on the dependence of the other parameters. For example, compared to the existing $R_{T}=O(\sqrt{T\ln^{d+2}T})$ regret under $k=k_{\mathrm{SE}}$, our regret upper bound $R_{T}=O(\sqrt{T\ln^{2}T})$ indeed avoids the dependence of $d$ in the logarithmic factor; however, under the joint limit of $d$ and $T$ ($d,T\rightarrow\infty$), it easily behaves super-linearly even under the slowly increasing $d$ (e.g., $d=\Theta(\ln\ln T)$) due to the hidden constants in the regret.
@@ 3.1 Intuitive Explanation of our Analysis
  • Case II: $I(\mathbf{X}_{T})\leq\gamma_{T}(\mathcal{X})$ is tight: Let us assume that $(\bm{x}_{t})$ is the same as the input sequence generated by the maximum variance reduction (MVR) algorithm (namely, $\forall t\in[T],\bm{x}_{t}\in\operatornamewithlimits{argmax}_{\bm{x}\in\mathcal{X}}\sigma(\bm{x};\mathbf{X}_{t-1})$[srinivas10gaussian, vakili2021optimal]. Then, from the discussion in Sections 2 and 5 in [srinivas10gaussian], we already know that $\gamma_{T}(\mathcal{X})\leq(1-1/e)^{-1}I(\mathbf{X}_{T})$. This suggests that $I(\mathbf{X}_{T})\leq\gamma_{T}(\mathcal{X})$ is tight up to a constant factor when $\mathbf{X}_{T}$ is realized by MVR.
@@
From Case I, we observe that $I(\mathbf{X}_{T})$ satisfies $\Theta(\ln T)\leq I(\mathbf{X}_{T})\leq\gamma_{T}(\mathcal{X})$ depending on $\mathbf{X}_{T}$. Furthermore, by comparing the input sequences in cases I and II, we expect that $I(\mathbf{X}_{T})$ becomes small if $\mathbf{X}_{T}$ concentrates around the neighborhood of $\bm{x}^{\ast}$, while $I(\mathbf{X}_{T})$ becomes large if $\mathbf{X}_{T}$ spreads over the entire input domain $\mathcal{X}$. Then, from the fact that the worst-case regret of GP-UCB increases sub-linearly with the speed of $O(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})})$, we can deduce that the input sequence $\mathbf{X}_{T}$ of GP-UCB will eventually concentrate around the maximizer $\bm{x}^{\ast}$ if $\bm{x}^{\ast}$ is unique and $\|f\|_{\infty}$ is not extremely small [^3] Specifically, if $T\|f\|_{\infty}\leq O(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})})$, we cannot make any claims about $\mathbf{X}_{T}$ based on the worst-case bound since any sequence $\mathbf{X}_{T}$ satisfies the worst-case bound without concentrating around maximizer. This is why our analysis technique does not improve the worst-case regret in the frequentist setting. Indeed, in the proof of the worst-case lower bound for the frequentist setting [scarlett2017lower], the existence of the function $f$ with $T\|f\|_{\infty}=O(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})})$ is guaranteed. . We provide an illustrative image in Figure 1. Our proof is designed so as to capture the above intuition that $I(\mathbf{X}_{T})$ could be improved from $\gamma_{T}(\mathcal{X})$ to $\Theta(\ln T)$ under “favorable” sample path $f$.
@@
Figure 1: The behavior of the realized input sequence $\mathbf{X}_{T}$ (left) and the corresponding information gain $I(\mathbf{X}_{T})$ (right) in the 1-dimensional BO problem with $\sigma^{2}=3$. The top left figure shows the objective function $f$ realized by GP under $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu=5/2$. The bottom left figure shows the histogram of the realized inputs: $(\bm{x}_{t})_{t\in[200]}$ with GP-UCB (blue) and MVR (orange) under $f$ in the top left figure. Furthermore, the right plot shows the corresponding information gain $I(\mathbf{X}_{t})$ under GP-UCB or MVR. We also plot $I(\mathbf{X}_{t})\coloneqq 0.5\ln(1+\sigma^{-2}t)$, corresponding to Case I described in Section 3.1. We can observe that the inputs selected by GP-UCB are concentrated around the maximizer from the left figure. Then, from the right figure, we also observe that the corresponding information gain increases more slowly than that of MVR, and behaves similarly to Case I on $t\geq 30$. More comprehensive empirical results are also provided in Appendix D.
@@ 3.2 Proof of Theorem 3
Let $\mathcal{A}$ be an event such that the two high-probability events of the original GP-UCB proof (described in the last paragraph in Section 2) and Lemma 2 with the confidence level $\delta_{\mathrm{GP}}$ simultaneously hold. Note that event $\mathcal{A}$ occurs with probability at least $1-\delta_{\mathrm{GP}}-\delta$ from the union bound. Therefore, it is enough to prove our upper bound under $\mathcal{A}$. To encode the high-level idea in the previous section, we need to capture the concentration behavior of the input sequence $\mathbf{X}_{T}$ around the maximizer $\bm{x}^{\ast}$. From this motivation, given some constant $\varepsilon>0$, we decompose the regret as $R_{T}=R_{T}^{(1)}(\varepsilon)+R_{T}^{(2)}(\varepsilon)$, where:
@@
We set $\mathcal{T}(\varepsilon)=\{t\in[T]\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})>\varepsilon\}$ and $\mathcal{T}^{c}(\varepsilon)=[T]\setminus\mathcal{T}(\varepsilon)$ in the above definition. A key observation is that, if we set sufficiently small $\varepsilon$ depending on the constants in Lemma 2, the inputs $(\bm{x}_{t})$ in $R_{T}^{(2)}(\varepsilon)$ (namely, inputs $(\bm{x}_{t})$ such that $f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon$ holds) are on the locally quadratic region around the maximizer $\bm{x}^{\ast}$ due to conditions 1 and 3 in Lemma 2. The formal descriptions are provided in Lemma 20 in Appendix C. This fact is originally leveraged in [scarlett2018tight] to analyze the successive elimination-based algorithm. In the analysis of GP-UCB, it enables us to analyze the behavior of the sub-input sequence $\{\bm{x}_{t}\mid f(\bm{x}^{\ast})-f(\bm{x}_{t})\leq\varepsilon\}$ through the regularity constant $c_{\mathrm{quad}}$. Below, we formally give the upper bound for $R_{T}^{(2)}(\varepsilon)$.
@@
Suppose $(\bm{x}_{t})_{t\in[T]}$ is the input query sequence realized by the GP-UCB algorithm. Furthermore, let $\overline{\gamma}_{t}$ is the upper bound of MIG $\gamma_{t}(\mathcal{X})$ such that $\overline{\gamma}_{t}/t$ is non-increasing on $\overline{T},\infty)$ with some $\overline{T}\in\mathbb{N}_{+}$ [^4] Namely, $\forall t\geq\overline{T},\forall\epsilon\geq 0,\overline{\gamma}_{t}/t\geq\overline{\gamma}_{t+\epsilon}/(t+\epsilon)$ and $\forall t\geq\overline{T},\gamma_{t}(\mathcal{X})\leq\overline{\gamma}_{t}$ hold for some $\overline{T}\in\mathbb{N}_{+}$.. Then, under event $\mathcal{A}$, we have
$R_{T}^{(2)}(\varepsilon)\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\max_{i\in[\overline{i}]}\sqrt{\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)},$
where $C=2/\ln(1+\sigma^{-2})$, $\overline{i}=\lfloor\log_{2}\frac{T}{\overline{T}}\rfloor+1$, $\eta_{i}=\frac{2\left(2\sqrt{C\beta_{T}(T/2^{i-1})\overline{\gamma}_{T/2^{i-1}}}+\frac{\pi^{2}}{6}\right)}{(T/2^{i-1})}$, and $\varepsilon=\min\{c_{\mathrm{gap}},c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}\}$.
@@
$R_{T}^{(2)}(\varepsilon)=\widetilde{O}\left(\max_{i}\sqrt{T\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)}\right).$ (12)
Note that $\eta_{i}$ is decreasing as the time index $T/2^{i-1}$ of MIG increases. In other words, the input domain $\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)$ of MIG shrinks as the time index $T/2^{i-1}$ increases. This property is beneficial for obtaining a tighter upper bound than that from the existing technique. For example, under $k=k_{\mathrm{Mat\acute{e}rn}}$ with $2\nu+d\leq\nu^{2}$, we can confirm that the dominant polynomial term in MIG is canceled out by the shrinking of the input domain in MIG. Namely, we can obtain the following result under $k=k_{\mathrm{Mat\acute{e}rn}}$:
$\max_{i}\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)=\widetilde{O}(1)~~(\mathrm{as}~~T\rightarrow\infty),$ (13)
@@
The full proof is given in Appendix A.2. The remaining interest is the upper bound of $R_{T}^{(1)}(\varepsilon)$. The definition of $R_{T}^{(1)}(\varepsilon)$ is the same as the lenient regret [cai2021lenient], which is known to be smaller than the original regret $R_{T}$ in GP-UCB. Although cai2021lenient studies the frequentist setting, their proof strategy is also applicable to the Bayesian setting as described in Section 3.4 in [cai2021lenient]. The following lemma provides the formal statement about the upper bound of $R_{T}^{(1)}(\varepsilon)$.
Lemma 6 (Upper bound of $R_{T}^{(1)}$, adaptation of the proof of Theorem 1 in [cai2021lenient]).